If alpha and wita are the zeroes of 2xsquare+5x-10 then what is the value of alpha*wita?
Answer by stanbon(75887)

If alpha and beta are the zeroes of 2xsquare+5x-10 then what is the value of alpha*beta?
-----
2x^2 + 5x - 10 = 0
----
Ans: [-b+sqrt(b^2-4ac)]/(2a) * [-b-sqrt(b^2-4ac)]/(2a)
=====
= [b^2 - (b^2-4ac)]/4a^2
====
= [b^2-b^2+4ac]/4a^2
---
= 4ac/4a^2
---
= c/a
----
Ans: alpha*beta = c/a = -10/2 = -5
------------------------
Cheers,
Stan H.
==================